In this talk we consider several new developments on strong connectivity questions in graphs. For a fixed multigraph H, possibly containing loops, with V (H) = {h_{1}, …, h_{k}}, we say a graph G is H-linked if for every choice of k vertices v_{1}, …, v_{k} in G, there exists a subdivision of H in G such that v_{i} represents h_{i} (for all i). Using a k-matching as the graph H, we obtain the well-known idea of a k-linked graph. An H-immersion in G is similar to H-linkage, except that the paths in G, playing the role of the edges of H, are only required to be edge disjoint. We determine minimum degree conditions for a graph G to contain an H-linkage. We further generalize these results to find conditions for a graph G to contain an H-immersion with a bounded number of vertex repetitions on any choice of k vertices.